# IPR-1: Interactive Physical Reasoner

### v1

@@ Abstract
Humans learn by observing, interacting with environments, and internalizing physics and causality. Here, we aim to ask whether an agent can similarly acquire human-like reasoning from interaction and keep improving with more experience. We study this in a Game-to-Unseen (G2U) setting, curating 1,000+ heterogeneous games with diverse physical and causal mechanisms, and evaluate at three human-like levels: Survival, Curiosity, Utility, from primitive intuition to goal-driven reasoning. Our analysis reveals complementary failures: VLM/VLA agents reason but lack look-ahead in interactive settings, while world models imagine but imitate visual patterns rather than analyze physics and causality. We therefore propose IPR (Interactive Physical Reasoner), using world-model rollouts to score and reinforce a VLM’s policy, and introduce PhysCode, a physics-centric action code aligning semantic intent with dynamics to provide a shared action space for prediction and reasoning. Pretrained on 1,000+ games, our IPR performs robustly on three levels, matches GPT-5 overall, and surpasses it on Curiosity. We find that performance improves with more training games and interaction steps, and that the model also zero-shot transfers to unseen games. These results support physics-centric interaction as a path to steadily improving physical reasoning. Our code will be publicly available.
footnotetext: Equal contribution.  Corresponding author. ∗∗Work conducted at Shanghai Jiao Tong University.
@@ 1 Introduction
We assume that, if an agent is exposed to diverse, interactive worlds and trained to capture shared physical and causal mechanisms, rather than domain-specific appearance or action interfaces, it would scale its physical reasoning ability reliably and transfer to new scenarios. This view aligns with prior reasoning works [Bengio+chapter2007, huang2023reasoninglargelanguagemodels, zhang2024stepbackrethinkingstages]. Large VLM/VLA models [wang2024jarvis, wang2025gametarspretrainedfoundationmodels, bu2025univlalearningacttaskcentric], pretrained on vast corpora, exhibit basic reasoning abilities and a certain degree of transfer; yet in interactive scenarios they lack forward prediction visually and often fail when precise anticipation of action consequences is required (e.g., timing, contact, momentum) (Fig. 2(b)). World models and other prediction-based methods [bruce2024genie, assran2025v, hafner2025trainingagentsinsidescalable] can imagine futures via differentiable latent dynamics and interactively optimize trajectories toward goal-aligned representations, but they tend to collapse into target-chasing imitation rather than genuine causal reasoning, failing in long-horizon and sparse-reward tasks. This naturally raises a question: can we blend the open-ended reasoning of VLMs with the predictive grounding of world models to support interactive reasoning in novel environments and yield competence that improves steadily with experience?
@@
(a) Three-level evaluation inspired by Maslow’s hierarchy of needs. We organize tasks into a pyramid of Survival, Curiosity, and Utility. Survival measures how long the agent can stay alive by avoiding risks; Curiosity measures how broadly it visits novel states; and Utility measures how well it achieves downstream goals. The three levels progress from physical intuition to goal-driven reasoning. Our IPR performs robustly across the entire pyramid.
(b) Motivating failure cases in control semantics, language grounding, and prediction. (1) Control conflict: the same key (e.g., UP) triggers different semantics across games (camera tilt up v.s. character move up), causing console aliasing. (2) Vision-language distortion: text-only actions cannot specify precise visual magnitudes (e.g., jump height/speed), leading to systematic amplitude errors. (3) Missing foresight: without imagination, the agent cannot anticipate upcoming hazards during interaction (e.g., spikes, moving enemies).
Figure 2: Overview: three-level evaluation pyramid (left) and failure cases of previous VLM-based model (right), motivating our IPR.
In this way, we propose IPR (Interactive Physical Reasoner): a paradigm where world model prediction reinforces a VLM policy to adapt its physical reasoning in interactive environments (Fig. 3). A natural obstacle arises when naively piping VLM outputs into a predictor: keyboard induces interface mismatches across games and language distorts visual details (Fig. 2(b)). We therefore introduce PhysCode, a physics-centric action code that fuses action semantics with visual dynamics into a compact discrete representation. Concretely, we follow a Genie [bruce2024genie] style discretization over fused features extracted from video frames, optical flow, and action semantics. Each code is encouraged to align with (i) domain-agnostic dynamical primitives (e.g., momentum change) and (ii) domain-specific visual affordances. Instead of issuing raw keys or free-form language, the reasoning policy outputs multiple PhysCode sequences, which are scored by the world model in the same latent space so that the best candidate would be executed and its imagined rewards are used to reinforce the policy.
To evaluate the paradigm at scale, we curate over 1,000 heterogeneous games spanning visual styles, control interfaces, and physical and causal mechanisms. Games form a low-cost, controllable testbed for physical reasoning: they afford rich interaction, physics closely resembling the real world, and effectively unlimited rollouts. We further organize evaluation into three levels inspired by Maslow’s hierarchy of needs [huitt2007maslow]: Survival, Curiosity, Utility, covering a spectrum from physical intuition to goal-directed reasoning (Fig. 2(a)). The result on three levels verifies two failure modes: reasoning-based VLM/VLA lack forward consequence prediction to explore (Curiosity), while prediction-based world models explore broadly yet fail at goal-driven tasks (Utility), which motivates our design.
@@
Across this suite, our IPR remains robust on all three levels, while RL-based and prediction-based baselines often collapse on one or more of them. Trained under the IPR paradigm, an 8B backbone matches GPT-5 overall and even surpasses it on curiosity. Moreover, competence scales with the number of training games and interaction steps (Fig. 5) and zero-shot transfers to novel environments, highlighting the potential of interactive learning for physical reasoning at scale.
@@ 2 Related Works
Research on action spaces spans hand-designed controls, language-based interfaces, and learned latent representations. Early embodied agents operated over environment-specific key bindings, torques, or joystick signals [mnih2015human, schulman2017proximal, haarnoja2018softactorcriticoffpolicymaximum, brockman2016openaigym], which offer precise control but entangle behavior with platform-specific layouts and hinder cross-domain transfer. A second line adopts language-based action spaces, issuing natural-language commands or tool calls [shridhar2020alfredbenchmarkinterpretinggrounded, fan2022minedojobuildingopenendedembodied, wang2024jarvis, wang2023voyager, ahn2022icanisay]; while language affords semantic generality, it abstracts away timing, force, and perception–action couplings, often leading to imprecise or under-grounded control [savva2019habitat, reed2022generalist]. A complementary direction learns latent action spaces directly from interaction data. Discrete or continuous latent codes—via VQ-VAE [oord2018neuraldiscreterepresentationlearning] or sequence models—have been explored for planning, control, and world models [bruce2024genie, chi2024diffusionpolicyvisuomotorpolicy, sharma2020emergentrealworldroboticskills, lynch2021languageconditionedimitationlearning, okada2021dreamingmodelbasedreinforcementlearning]. Recent VLM/VLA systems integrate such latent tokens into large multimodal models [sapkota2025vision, huang2024embodiedgeneralistagent3d], but these codes often remain entangled across domains and lack mechanisms to capture shared physical principles versus environment-specific affordances. Our work addresses this gap by learning a physics-centric latent action space that captures reusable dynamical patterns across games, instead of binding actions to domain-specific visuals and control layouts.
@@
Research on game-playing agents has largely followed three threads. RL-based agents, from DQN and PPO/SAC to large-scale systems like AlphaStar and OpenAI Five [mnih2013playing, vanhasselt2015deepreinforcementlearningdouble, schulman2017proximal, haarnoja2018softactorcriticoffpolicymaximum, Vinyals2019GrandmasterLI, openai2019dota2largescale], learn policies directly from pixels and rewards and achieve strong title-specific performance, but remain sample-inefficient, brittle to interface changes, and struggle with long-horizon credit assignment and cross-game transfer. Prediction-based (world-model) agents such as World Models, PlaNet, the Dreamer family, and Genie [https://doi.org/10.5281/zenodo.1207631, hafner2019learninglatentdynamicsplanning, hafner2020dreamcontrollearningbehaviors, hafner2023mastering, bruce2024genie] first learn latent dynamics and then plan or optimize in imagination, improving exploration and sparse-reward learning, yet degrade when learned dynamics or action semantics drift from the test environment and typically optimize task or pixel losses rather than reasoning quality. VLM/VLA-based agents like Gato, RT-2, Voyager, MineDojo, and recent VLA frameworks [reed2022generalist, brohan2023rt2visionlanguageactionmodelstransfer, wang2023voyager, fan2022minedojobuildingopenendedembodied] cast acting as sequence modeling over images, text, and actions and excel at zero-shot instruction following, but rely heavily on static corpora, heuristic wrappers, and weakly grounded forward prediction. Our IPR paradigm aims to inherit the strengths of these lines by using a physics-centric latent action space where a world model provides imagination-based value estimates and a reasoning VLM policy is reinforced through interactive experience in the same latent space.
@@
Interactive environments have long served as testbeds for learning control, exploration, and generalization: Atari/ALE provided dense stepwise rewards for RL training and evaluation [mnih2013playing, bellemare2013arcade], while later platforms such as Minecraft, VizDoom, and StarCraft introduced long-horizon goals, partial observability, and sparse rewards [kempka2016vizdoom, vinyals2017starcraft, fan2022minedojobuildingopenendedembodied, wang2023voyager]. With the rise of VLM/VLA agents, web-based benchmarks and browser environments have been proposed to test generalization to novel tasks and interfaces [yuan2023rl, qi2024webrl]. Following this line, we evaluate agents on a diverse suite of games and adopt simple game-agnostic metrics grouped into three levels—survival, curiosity, and utility—to provide their preformance from physical intuition to reasoning and their scaling with experience.
@@ 5 Experiments
5.1 Setup: Datasets, Tasks, and Metrics
Sources.
Diversity axes.
@@ 5.1 Setup: Datasets, Tasks, and Metrics
Inspired by Maslow’s hierarchy of needs [huitt2007maslow], we treat gameplay as a three-level progression: Survival $\rightarrow$ Curiosity $\rightarrow$ Utility (Fig. 2(a)), from intuition to reasoning.
@@
Table 2: Comprehensive comparison across [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. Obj: training objective; Mean: normalized score; Avg. Rank: normalized average rank (1/(n1)1/(n{-}1), higher is better); Ratio@Top-3(%): percentage of games where the method is in the top-3.
Methods [Uncaptioned image]Survival [Uncaptioned image]Curiosity [Uncaptioned image]Utility
Mean Avg. Rank Ratio@Top-3(%) Mean Avg. Rank Ratio@Top-3(%) Mean Avg. Rank Ratio@Top-3(%)
Control Group
Random 0.541 0.450 3.5 5.254 0.363 5.8 0.000 0.559 1.1
Human 0.764 0.760 41.9 5.428 0.480 11.7 1.000 0.874 76.1
Imitation Learning (IL) Group
ACT-BC 0.541 0.474 5.8 6.896 0.476 13.0 0.092 0.487 5.7
Qwen3-VL-8B-BC 0.578 0.578 7.0 6.945 0.502 11.8 0.129 0.560 3.4
Reinforcement Learning (RL) Group
PPO@survival 0.571 0.581 16.3 5.453 0.415 2.3 0.259 0.636 14.8
PPO@curiosity 0.562 0.566 11.6 5.344 0.373 3.5 0.285 0.659 19.3
PPO@utility 0.554 0.609 17.4 5.419 0.397 2.3 0.268 0.637 18.2
DQN@survival 0.538 0.493 15.1 7.058 0.588 10.6 0.147 0.518 8.0
DQN@curiosity 0.562 0.531 18.6 6.841 0.578 12.8 0.159 0.521 8.4
DQN@utility 0.556 0.515 18.6 6.450 0.563 9.3 0.156 0.536 9.1
World Model Group
DreamerV3@survival 0.521 0.495 8.1 7.737 0.526 14.0 0.114 0.446 6.8
DreamerV3@curiosity 0.538 0.466 8.1 7.843 0.608 20.9 0.138 0.449 11.5
DreamerV3@utility 0.538 0.481 9.3 7.336 0.530 17.6 0.139 0.446 8.7
V-JEPA2@survival 0.531 0.394 5.8 7.815 0.469 14.0 0.065 0.285 4.0
V-JEPA2@curiosity 0.516 0.343 0.0 8.463 0.610 22.1 0.046 0.246 4.5
V-JEPA2@utility 0.526 0.369 1.2 7.777 0.412 15.1 0.051 0.293 4.5
GenieRedux@survival 0.539 0.435 3.5 7.937 0.573 17.4 0.090 0.334 4.0
GenieRedux@curiosity 0.534 0.421 3.5 8.390 0.675 20.9 0.084 0.293 3.4
GenieRedux@utility 0.540 0.445 4.7 8.067 0.573 17.4 0.085 0.349 4.5
Multimodal Large Language Model (MLLM) Group
GPT-4o@survival 0.575 0.605 14.0 4.954 0.390 2.2 0.156 0.543 7.0
GPT-4o@curiosity 0.521 0.429 7.0 5.800 0.518 4.7 0.124 0.474 4.5
GPT-4o@utility 0.522 0.474 8.1 5.676 0.465 2.3 0.162 0.549 8.0
GPT-5@survival 0.619 0.630 25.6 5.181 0.358 2.3 0.199 0.593 18.2
GPT-5@curiosity 0.529 0.488 5.8 5.621 0.448 5.8 0.148 0.515 5.7
GPT-5@utility 0.536 0.460 5.8 5.329 0.427 2.3 0.206 0.606 13.6
Qwen3-VL-A30B@survival 0.544 0.514 5.8 5.690 0.522 5.8 0.149 0.539 6.8
Qwen3-VL-A30B@curiosity 0.518 0.450 7.0 7.113 0.600 11.7 0.133 0.500 4.5
Qwen3-VL-A30B@utility 0.545 0.504 9.3 6.214 0.443 4.7 0.146 0.552 6.8
Interactive Physical Reasoner
IPR (8B) 0.589 0.517 10.4 7.874 0.584 15.6 0.178 0.604 8.9
(ranking) (2/26) (8/26) (9/26) (5/26) (6/26) (7/26) (7/26) (5/26) (9/26)
Key Takeaways across [Uncaptioned image]Survival, [Uncaptioned image]Curiosity, and [Uncaptioned image]Utility Prediction-based Methods (WM). Strong at [Uncaptioned image], but weaker at [Uncaptioned image] and [Uncaptioned image]. Trained on broad exploratory trajectories, latent rollouts broaden coverage and reveal dynamics, but tend to imitate visually-alike futures rather than reliably pursue goals. So prediction is useful as a look-ahead prior for risk and candidate actions. RL-based Methods (PPO, DQN). Strong at [Uncaptioned image]and [Uncaptioned image] when rewards are well-shaped, but weaker on [Uncaptioned image] and tasks without explicit goals. Reward gradients enable effective credit assignment under the right signal, yet sparsity and partial observability induce instability and interface overfitting—so RL works best as an optimization method. Experience-based Methods (Behavior Cloning). Strong at human-like [Uncaptioned image], but weaker on [Uncaptioned image] and [Uncaptioned image]. Deliberately imitate human trajectories and thus excel at low-risk survival, but struggle once tasks require precise control or exploration, and their performance depends strongly on the coverage and quality of the demonstrations. Reasoning-based Pretrained VLMs. Strong at goal-conditioned [Uncaptioned image]and [Uncaptioned image]; weaker on [Uncaptioned image]. They excel at instruction-driven reasoning but cannot predict consequences in the visual state space, so they work best as high-level reasoners that need auxiliary prediction modules for outcome-aware decisions. Interactive Physical Reasoning (Ours). Robust across [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]. We combine the strengths of all three paradigms: VLMs provide goal-driven causal reasoning, the world model supplies rollout prediction, and RL optimizes decisions using imagined rewards, yielding consistently strong performance across all three levels.

### v2

@@ Abstract
Humans learn by observing, interacting with environments, and internalizing physics and causality. Here, we aim to ask whether an agent can similarly acquire human-like reasoning from interaction and keep improving with more experience. To study this, we introduce a Game-to-Unseen (G2U) benchmark of 1,000+ heterogeneous games that exhibit significant visual domain gaps. Existing approaches, including VLMs and world models, struggle to capture underlying physics and causality since they are not focused on core mechanisms and overfit to visual details. VLM/VLA agents reason but lack look-ahead in interactive settings, while world models imagine but imitate visual patterns rather than analyze physics and causality. We therefore propose IPR (Interactive Physical Reasoner), using world-model rollouts to score and reinforce a VLM’s policy, and introduce PhysCode, a physics-centric action code aligning semantic intent with dynamics to provide a shared action space for prediction and reasoning. Pretrained on 1,000+ games, our IPR performs robustly on levels from primitive intuition to goal-driven reasoning, and even surpasses GPT-5 overall. We find that performance improves with more training games and interaction steps, and that the model also zero-shot transfers to unseen games. These results support physics-centric interaction as a path to steadily improving physical reasoning. Further demos and project details can be found at https://mybearyzhang.github.io/ipr-1.
footnotetext: Equal contribution.  Corresponding author. ∗∗Conducted during an internship at Shanghai Jiao Tong University.
@@ 1 Introduction
Figure 2: Three-level evaluation inspired by Maslow’s hierarchy of needs. We organize tasks into a pyramid of Survival, Curiosity, and Utility. Survival measures how long the agent can stay alive by avoiding risks. Curiosity measures how broadly it visits novel states; and Utility measures how well it achieves downstream goals. The three levels progress from physical intuition to goal-driven reasoning. Our IPR performs robustly across the entire pyramid.
We assume that, if an agent is exposed to diverse, interactive worlds and trained to distill shared physical and causal mechanisms, rather than domain-specific appearance or action interfaces, it would scale its physical reasoning ability reliably and transfer to new scenarios. This view resonates with prior reasoning works [Bengio+chapter2007, huang2023reasoninglargelanguagemodels, zhang2024stepbackrethinkingstages]. Pretrained VLMs [hu2024survey, wang2025gametarspretrainedfoundationmodels], despite broad semantic priors from large-scale pretraining, mainly acquire static, pattern-matching behaviors as open-loop planning; SFT strengthens text-based declarative reasoning but not the predictive grounding required for interactive physical tasks. Behavior-cloned VLAs [wang2024jarvis, bu2025univlalearningacttaskcentric] are upper-bounded by the quality of demonstrations; relying on rote memorization of trajectories, they lack adaptability and prone to failure under environmental variations. Model-based approaches [wiener2019cybernetics, moerland2023model], including control theory and model-based RL, ensure stability with accurate dynamics but struggle in complex environments where faithful modeling is infeasible. Model-free RL [yu2020meta, osband2016deep] avoids explicit modeling but demands massive samples and dense rewards. It often overfits to task-specific shortcuts rather than causal mechanisms, hindering transfer to complex tasks. Recent world models and other prediction-based approaches [bruce2024genie, assran2025v, hafner2025trainingagentsinsidescalable] scale effectively by learning latent dynamics and enabling agents to imagine futures. They can optimize actions interactively toward goal-aligned representations, but in practice they often collapse into short-horizon target chasing or imitation of surface correlations, lacking robust causal reasoning and suffering from compounding errors in complex environments.
Collectively, these limitations highlight a fundamental gap: while existing paradigms exhibit partial success, they tend to overfit to superficial visual details rather than capturing the underlying physical and causal mechanisms. Appropriately approximating these invariant dynamics, which are pivotal for robust transfer across interactive environments, requires leveraging diverse domains to disentangle core mechanisms from visual appearance. Reinforcement Learning (RL) excels at optimizing by interaction but relies on sparse, task-entangled signals that hinder generalization; Generative World Models capture dynamics but often over-models the full sensory space; and VLMs, despite offering rich semantic priors, lack the predictive grounding required for precise physical consistency. This motivates a “blended” perspective: instead of committing entirely to exploration (RL), full-scene prediction (world models), or static priors (VLMs), we should reconsider the ratio to absorb these components. To operationalize this, our approach aligns with the Latent World Model paradigm [assran2025v], which serves as the structural backbone to integrate these strengths. Specifically, a scalable reasoner should (i) model only the essential latent dynamics necessary for anticipating consequences—discarding high-fidelity pixel reconstruction; (ii) interact with raw multimodal signals through a policy enriched by VLM-based semantic priors; and (iii) reinforce this policy using predictive feedback that reflects physical feasibility. By shifting the prediction target from raw observations to abstract representations, the system filters out task-irrelevant perceptual noise, allowing the agent to capture the “essence” of physical and causal mechanisms rather than the “appearance” of the world.
In this way, we propose IPR (Interactive Physical Reasoner), a paradigm where world model prediction reinforces a VLM policy to adapt its physical reasoning in interactive environments (Fig. 5). To evaluate this paradigm at scale, we curate over 1,000 heterogeneous games spanning diverse visual styles, control interfaces, physics configurations, and causal structures. Games provide an ideal testbed for physical reasoning: they offer rich interaction, realistic physics, and effectively unlimited rollouts at low cost. Crucially, their heterogeneous visual appearances introduce substantial domain gaps that typically break traditional agents trained environment-by-environment. For IPR, however, these diverse worlds share the same underlying physical and causal principles, allowing it to learn a representation focused enough to transfer across radically different domains.
@@
We further organize evaluation into three levels inspired by Maslow’s hierarchy of needs [huitt2007maslow]: Survival, Curiosity, Utility, covering a spectrum from physical intuition to goal-directed reasoning (Fig. 2). The result on three levels verifies two failure modes: reasoning-based VLM/VLA lack forward consequence prediction to explore (Curiosity), while prediction-based world models explore broadly yet fail at goal-driven tasks (Utility). Across the full suite, our IPR remains robust on all three levels, while RL-based and prediction-based baselines often collapse on one or more of them. With an 8B backbone, IPR even surpasses GPT-5 overall. Moreover, competence scales with the number of training games and interaction steps (Fig. 7) and zero-shot transfers to novel environments, highlighting the potential of interactive learning for physical reasoning at scale. We will further extend this paradigm to real-world interactive environments and perform on robotic tasks.
@@ 2 Related Works
Research on action spaces spans hand-designed controls, language-based interfaces, and learned latent representations. Early embodied agents operated over environment-specific key bindings, torques, or joystick signals [mnih2015human, schulman2017proximal, haarnoja2018softactorcriticoffpolicymaximum, brockman2016openaigym], which offer precise control but entangle behavior with platform-specific layouts and hinder cross-domain transfer. A second line adopts language-based action spaces, issuing natural-language commands or tool calls [shridhar2020alfredbenchmarkinterpretinggrounded, fan2022minedojobuildingopenendedembodied, wang2024jarvis, wang2023voyager, ahn2022icanisay]; while language affords semantic generality, it abstracts away timing, force, and perception–action couplings, often leading to imprecise or under-grounded control [savva2019habitat, reed2022generalist]. A complementary direction learns latent action spaces directly from interaction data. Discrete or continuous latent codes, via VQ-VAE [oord2018neuraldiscreterepresentationlearning] or sequence models, have been explored for planning, control, and world models [bruce2024genie, chi2024diffusionpolicyvisuomotorpolicy, sharma2020emergentrealworldroboticskills, lynch2021languageconditionedimitationlearning, okada2021dreamingmodelbasedreinforcementlearning]. Recent VLM/VLA systems integrate such latent tokens into large multimodal models [sapkota2025vision, huang2024embodiedgeneralistagent3d], but these codes often remain entangled across domains and lack mechanisms to capture shared physical principles versus environment-specific affordances. Our work addresses this gap by learning a physics-centric latent action space that captures reusable dynamical patterns across games, instead of binding actions to domain-specific visuals and control layouts. Fig. 4 shows that different worlds share some semantic actions, validating the design of shared latent action space.
Figure 3: Motivating failure cases in control semantics, language grounding, and prediction. (1) Control conflict: the same key (e.g., UP) triggers different semantics across games (camera tilt up v.s. character move up), causing console aliasing. (2) Vision-language distortion: text-only actions cannot specify precise visual magnitudes (e.g., jump height/speed), leading to systematic amplitude errors. (3) Missing foresight: without imagination, the agent cannot anticipate upcoming hazards during interaction (e.g., spikes, moving enemies).
Figure 4: Word cloud of action semantics across thousands of game worlds. These shared semantics provide the structural foundation for cross-domain transfer. Actions highlighted in red represent those shared with general robotic operations, while the size of each word reflects its frequency in our data recipe.
@@
Research on game-playing agents has largely followed three threads. RL-based agents, from DQN and PPO/SAC to large-scale systems like AlphaStar and OpenAI Five [mnih2013playing, vanhasselt2015deepreinforcementlearningdouble, schulman2017proximal, haarnoja2018softactorcriticoffpolicymaximum, Vinyals2019GrandmasterLI, openai2019dota2largescale], learn policies directly from pixels and rewards and achieve strong title-specific performance, but remain sample-inefficient, brittle to interface changes, and struggle with long-horizon credit assignment and cross-game transfer. Prediction-based (world-model) agents such as World Models, PlaNet, the Dreamer family, and Genie [https://doi.org/10.5281/zenodo.1207631, hafner2019learninglatentdynamicsplanning, hafner2020dreamcontrollearningbehaviors, hafner2023mastering, bruce2024genie] first learn latent dynamics and then plan or optimize in imagination, improving exploration and sparse-reward learning, yet degrade when learned dynamics or action semantics drift from the test environment and typically optimize task or pixel losses rather than reasoning quality. VLM/VLA-based agents like Gato, RT-2, Voyager, MineDojo, and recent VLA frameworks [reed2022generalist, brohan2023rt2visionlanguageactionmodelstransfer, wang2023voyager, fan2022minedojobuildingopenendedembodied] cast acting as sequence modeling over images, text, and actions and excel at zero-shot instruction following, but rely heavily on static corpora, heuristic wrappers, and weakly grounded forward prediction (Fig. 3). Our IPR paradigm aims to inherit the strengths of these lines by using a physics-centric latent action space where a world model provides imagination-based value estimates and a reasoning VLM policy is reinforced through interactive experience in the same latent space.
@@
Interactive environments have long served as testbeds for learning control, exploration, and generalization: Atari/ALE provided dense stepwise rewards for RL training and evaluation [mnih2013playing, bellemare2013arcade], while later platforms such as Minecraft, VizDoom, and StarCraft introduced long-horizon goals, partial observability, and sparse rewards [kempka2016vizdoom, vinyals2017starcraft, fan2022minedojobuildingopenendedembodied, wang2023voyager]. With the rise of VLM/VLA agents, web-based benchmarks and browser environments have been proposed to test generalization to novel tasks and interfaces [yuan2023rl, qi2024webrl]. Following this line, we evaluate agents on a diverse suite of games and adopt simple game-agnostic metrics grouped into three levels: survival, curiosity, and utility, to provide their preformance from physical intuition to reasoning and their scaling with experience.
@@ 5 Experiments
5.1 Setup: Datasets, Tasks, and Metrics
Sources.
Diversity axes.
@@ 5.1 Setup: Datasets, Tasks, and Metrics
Inspired by Maslow’s hierarchy of needs [huitt2007maslow], we treat gameplay as a three-level progression: Survival $\rightarrow$ Curiosity $\rightarrow$ Utility (Fig. 2), from intuition to reasoning.